a piece of chromium metal with a mass of 24.26g is heated in boiling water to 98.3 C and then dropped into a coffee-cup calorimeter containing 82.3g of water at 23.3 C. When thermal equilibrium is reached the final temperature is 25.6 C. Calculate the specific heat capacity of chromium


1
Expert's answer
2021-10-08T02:07:01-0400

1) This problem can be summarized thusly:

-qlost by chromium = qgained by water

 

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J oC-1 g-1)

Tf = final temperature (oC)

Ti = initial temperature (oC)

 

The specific heat capacity of water is 4.20 J oC-1 g-1


2) Thus:

qlost by chromium = (24.26 g) × Cchromium × (25.6 - 98.3)°C

qlost by chromium = -1763.702 × Cchromium

 

qgained by water = (82.3 g) × (4.20 J oC-1 g-1) × (25.6 - 23.3)°C

qgained by water = 795.018 J


3) Therefore,

-qlost by chromium = qgained by water

(1763.702 oC g) × Cchromium = (795.018 J)

Cchromium = (795.018 J) / (1763.702 oC g) = 0.45 J oC-1 g-1

Cchromium = 0.45 J oC-1 g-1


Answer: The specific heat capacity of chromium is 0.45 J oC-1 g-1